# Next-Generation Reservoir Computing for Dynamical Inference

### v2

@@ I Introduction
Reservoir computing (RC) has emerged as a powerful and versatile machine learning framework for processing temporal data, particularly in tasks involving time series prediction and the inference of latent variables in dynamical systems [3, 4, 5]. As Lu, Hunt, and Ott [1] describe, “Reservoir computing is a machine-learning approach that has demonstrated success at a variety of tasks, including time series prediction and inferring unmeasured variables of a dynamical system from measured variables”. In its classical form, RC leverages a high-dimensional, nonlinear dynamical system — referred to as the reservoir — which is driven by an input signal. The reservoir’s response is then post-processed with a linear map to obtain the desired output, commonly referred to as readout. For a general overview of the theory and state of the art, see also Refs. [6, 7].
@@
NGRC departs from the traditional architecture by eliminating the explicit reservoir altogether. Instead, it constructs a high-dimensional feature space of dimension, $M$, (corresponding to an implicit reservoir) through a deterministic projection of the input signal. This results in a simple, fully transparent mapping from one time step to the next, where the system state is entirely known and controllable. Letting $u(t)$ denote the input and $r(t)$ the projected state, the evolution becomes a direct function $r(t+1)=F(r(t),u(t))$, enabling precise interventions and measurements. This is particularly advantageous in applications requiring system perturbation and response analysis, such as the computation of phase response curves [18, 19, 20, 21, 22]. See Sec. 1(b) and Fig. 1 for a more detailed comparison of the two methods.
@@ II.0.1 Traditional Reservoir Computing (RC)
In reservoir computing, the equations governing the reservoir dynamics and output are used during both the training and predicting phases. During the training phase, the reservoir state $\mathbf{r}(t)$ evolves according to the input signal $\mathbf{u}(t)$, and the output $\mathbf{u}(t)=\mathbf{W}_{\text{out}}\cdot\mathbf{r}(t)$ is used for learning the mapping represented by the weight matrix $\mathbf{W}_{\text{out}}$. In the prediction phase, the trained output layer $\mathbf{W}_{\text{out}}$ is applied to the reservoir state to determine the new input signal $\mathbf{u}(t)$, which then in turn continues drive the evolution of the reservoir. (We note that some literature distinguishes variables for the training and prediction phases, but for our purposes we maintain the same notation for $\mathbf{r}(t)$ and $\mathbf{u}(t)$, regardless of whether they represent values for training or prediction phases.) In this context, readout weights are parameters optimized during the training phase, whereas the projection dimension $M$ and other process defining quantities often are referred to as hyperparameters, are set before training and held fixed during prediction.
@@ IV.1.2 Rössler system
Figure 9 compares the true Rössler attractor with the attractor reconstructed from the reservoir model using delay embedding. Again, the original and reconstruced attractors match very well.
@@ IV.2 Reconstruction of bifurcation diagram
We consider the Rössler system (12) and add a variable control parameter $\eta$ to the right-hand-side of the $y$ equation. The reservoir then observes the resulting system states and learns to reproduce the system’s behaviour for different $\eta$. The variation of $\eta$ allows us to probe the bifurcation structure of the system. In the following, we first use several data sections corresponding to different constant $\eta$ values to reconstruct the bifurcation diagram, and secondly, consider a time-dependent, stochastic $\eta(t)$ to mimic less controlled real world situations.
@@ V Discussion
Data sharing is not applicable to this article as no new data were created or analyzed in this study. For reproducibility, all code used in this work is publicly available at https://github.com/rok-cestnik/next-gen-res-comp.
@@ Acknowledgements.
We thankfully acknowledge useful discussions with Andrew Flynn and Herbert Jaeger. We acknowledge financial support from the Crafoord Foundation (project no. 20240689).

### v3

@@ I Introduction
Reservoir computing (RC) has emerged as a powerful and versatile machine learning framework for processing temporal data, particularly in tasks involving time series prediction and the inference of latent variables in dynamical systems [3, 4, 5]. As Lu et al. [1] describe, “Reservoir computing is a machine-learning approach that has demonstrated success at a variety of tasks, including time series prediction and inferring unmeasured variables of a dynamical system from measured variables”. In its classical form, RC leverages a high-dimensional, nonlinear dynamical system — referred to as the reservoir — which is driven by an input signal. The reservoir’s response is then post-processed with a linear map to obtain the desired output, commonly referred to as readout. For a general overview of the theory and state of the art, see also Refs. [6, 7].
@@
NGRC departs from the traditional architecture by eliminating the explicit reservoir altogether. Instead, it constructs a high-dimensional feature space of dimension, $M$, (corresponding to an implicit reservoir) through a deterministic projection of the input signal. This results in a simple, fully transparent mapping from one time step to the next, where the system state is entirely known and controllable. Letting $u(t)$ denote the input and $r(t)$ the projected state, the evolution becomes a direct function $r(t+1)=F(r(t),u(t))$, enabling precise interventions and measurements. This is particularly advantageous in applications requiring system perturbation and response analysis, such as the computation of phase response curves [18, 19, 20, 21, 22]. See Sec. II and Fig. 1 for a more detailed comparison of the two methods.
@@ II.0.1 Traditional Reservoir Computing (RC)
In reservoir computing, the equations governing the reservoir dynamics and output are used during both the training and predicting phases. During the training phase, the reservoir state $\mathbf{r}(t)$ evolves according to the input signal $\mathbf{u}(t)$, and the output $\mathbf{u}(t)=\mathbf{W}_{\text{out}}\cdot\mathbf{r}(t)$ is used for learning the mapping represented by the weight matrix $\mathbf{W}_{\text{out}}$. In the prediction phase, the trained output layer $\mathbf{W}_{\text{out}}$ is applied to the reservoir state to determine the new input signal $\mathbf{u}(t)$, which then in turn continues drive the evolution of the reservoir. (We note that some literature distinguishes variables for the training and prediction phases, but for our purposes we maintain the same notation for $\mathbf{r}(t)$ and $\mathbf{u}(t)$, regardless of whether they represent values for training or prediction phases.) In this context, readout weights are parameters optimized during the training phase, whereas the projection dimension $M$ and other process defining quantities often are referred to as hyperparameters, and are set before training and held fixed during prediction.
@@ IV.1.2 Rössler system
Figure 9 compares the true Rössler attractor with the attractor reconstructed from the reservoir model using delay embedding. Again, the original and reconstructed attractors match very well.
@@ IV.2 Reconstruction of bifurcation diagram
We consider the Rössler system (12) and add a variable control parameter $\eta$ to the right-hand-side of the $y$ equation. The reservoir then observes the resulting system states and learns to reproduce the system’s behavior for different $\eta$. The variation of $\eta$ allows us to probe the bifurcation structure of the system. In the following, we first use several data sections corresponding to different constant $\eta$ values to reconstruct the bifurcation diagram, and secondly, consider a time-dependent, stochastic $\eta(t)$ to mimic less controlled real world situations.
@@ V Discussion
Data sharing is not applicable to this article as no new data were created or analyzed in this study. For reproducibility, all code used in this work is publicly available at https://github.com/rok-cestnik/next-gen-res-comp, Ref. [47].
@@ Acknowledgements.
We thankfully acknowledge useful discussions with Irmantas Ratas, Andrew Flynn, Edmilson Roque dos Santos and Herbert Jaeger. We acknowledge financial support from the Crafoord Foundation (project no. 20240689).